### v1

@@ 7.1 Jupyter Notebooks Code Quality
One of the activities that programmers engage in during the code comprehension phase of programming is executing different parts of the code and monitoring state variables. In a study by Wang et al. wang2020assessing , 10,000 GitHub repositories containing Jupyter notebooks were randomly cloned. They found that 73% of the repositories were not reproducible, which can have a negative impact on code understanding.
@@
While the ability to execute code cells in different orders can help data scientists try out different settings and find the best one, it also presents a challenge for the reproducibility of results. The order of code cells in a notebook may not match the order in which the original developer executed them. To address this issue, Wang et al. developed a tool that can automatically find the possible orders that satisfy the dependencies between code cells wang2020assessing .
Wang et al. wang2020better used the PEP8 checker to analyze the code quality in Jupyter notebooks shared in two Kaggle competitions. The PEP8 checker validates whether Python code follows Python coding style best practices, such as line length and spacing between operators. They observed that 36.26% of lines of code in Jupyter notebooks led to PEP8 checker errors, which may negatively affect code understanding. In contrast, when the PEP checker was applied to another dataset of Python scripts, only 13% of Python lines were erroneous. This study clearly demonstrates that the quality of Jupyter notebook code is significantly lower than that of regular Python scripts based on PEP checker metrics. Grotov et al. grotov2022large conducted a comparison between Python code written in Jupyter notebooks and traditional Python scripts, examining them from two perspectives: structural and stylistic. In terms of structure, they discovered several distinct aspects in which notebook code differs from its traditional counterpart. Notebooks exhibited a unique pattern in the utilization of functions, and the complexity metrics indicated that they contain code that is structurally simpler but more intertwined. This could potentially decrease their comprehensibility. In terms of stylistic differences, the disparities were more pronounced, with notebooks having 1.4 times more stylistic errors compared to scripts.
@@
The studies mentioned attempted to measure code quality using various features derived from Jupyter notebooks, but none of them measured the extent to which each feature can affect CU. Some of these studies only analyzed a small number of notebooks, which may not be representative of the broader population of Jupyter notebooks on GitHub. There is currently no widely accepted set of metrics for evaluating Jupyter notebook code quality. Different studies may use different metrics, making it difficult to compare results across studies. Our study is different and complements these studies by applying a machine learning model to determine if each metric is correlated with CU. We also report which features have a significant impact on code comprehension and which features have an insignificant effect.
7.2 Measuring Code Understandability Based on Opinion
@@
Scalabrino et al. scalabrino2019automatically conducted an extensive evaluation of 121 existing, as well as new, code-related, documentation-related, and developer-related metrics. They attempted to correlate each metric with understandability and build models that combine metrics to assess understandability. To do this, they used 444 human evaluations from 63 developers. One of the criteria they used to measure CU is called Perceived Binary Understandability (PBU), which is based on developers’ opinions. This is a binary categorical variable that is true if a developer perceives that they understood a given code, and false otherwise. Finally, they obtained a bold negative result: none of the 121 experimented metrics is able to capture CU. As in the previous article, the methodology of this study is based on collecting opinions from a limited number of developers. The research scope of this article was limited to Java-based software. PBU criteria were also used in scalabrino2019automatically for CU measurement.
@@
The main challenges posed by the approaches presented in these studies are that they solely rely on user votes on software repositories. These metrics may be affected by factors beyond code comprehensibility, including ease of use, algorithm accuracy, popularity, and marketing considerations.
@@ 8.1 Internal Validity
To ensure notebooks garnered enough user engagement to gather varied opinions, we set a baseline of 500 views. Notebooks falling short of this mark were excluded from our analysis.
@@ 8.3 Construct Validity
Construct validity is defined as an assessment of whether the theoretical constructs are being accurately understood and measured. This involves testing whether the interpretation of the constructs aligns with their intended meanings easterbrook2008selecting .
To ensure construct validity, each feature in our dataset is extracted based on the procedure explained in Section 3. Other members review the implementation by each author, and the errors are communicated to the developers. Even though we used a peer-review process to minimize the risk of failures in our implementations, It is still possible for some logical errors to remain unseen.
@@ 8.4 Replicability
We presented the experimental setup through multiple Jupyter notebooks. Moreover, the dataset and source code for training and evaluation of our models are publicly available online, allowing fellow researchers to replicate our results.
@@ 9 Conclusions and Future Work
Data Availability Statement.
@@
Moreover, our service, which accepts a notebook and returns 33 code metrics and predict its CU level, is accessible at http://46.249. 101.150/docs.
Conflict of Interest.
@@
Appendix A Distribution of Comments
@@
Appendix B Performance Tier Calculation

### v2

@@ 8.1 Jupyter Notebooks Code Quality
One of the activities that programmers engage in during the code comprehension phase of programming is executing different parts of the code and monitoring state variables. In a study by Wang et al. wang2020assessing , 10,000 GitHub repositories containing Jupyter notebooks were randomly cloned. They found that 73% of the repositories were not reproducible, which can have a negative impact on CU.
@@
While the ability to execute code cells in different orders can help data scientists try out different settings and find the best one, it also presents a challenge to the reproducibility of results. The order of code cells in a notebook may not match the order in which the original developer executed them. To address this issue, Wang et al. developed a tool that can automatically find the possible orders that satisfy the dependencies between code cells wang2020assessing .
Wang et al. wang2020better used the PEP8 checker to analyze the code quality in Jupyter notebooks shared in two Kaggle competitions. The PEP8 checker validates whether Python code follows Python coding style best practices, such as line length and spacing between operators. They observed that 36.26% of lines of code in Jupyter notebooks led to PEP8 checker errors, which may negatively affect CU. In contrast, when the PEP checker was applied to another dataset of Python scripts, only 13% of Python lines were erroneous. Our study clearly demonstrates that the quality of Jupyter notebook code is significantly lower than that of regular Python scripts based on PEP checker metrics. Grotov et al. grotov2022large conducted a comparison between Python code written in Jupyter notebooks and traditional Python scripts, examining them from two perspectives: structural and stylistic. In terms of structure, they discovered several distinct aspects in which notebook code differs from its traditional counterpart. Notebooks exhibited a unique pattern in the utilization of functions, and the complexity metrics indicated that they contain code that is structurally simpler but more intertwined. This could potentially decrease their comprehensibility. In terms of stylistic differences, the disparities were more pronounced, with notebooks having 1.4 times more stylistic errors compared to scripts.
@@
Brown et al. brown2023facilitating explore methods for facilitating dependency exploration in computational notebooks by focusing on tools and visualization techniques that help users understand cell dependencies, thereby enhancing reproducibility and comprehensibility of notebook workflows. Their approach primarily addresses the structural challenges of navigating inter-cell dependencies, offering visualization aids like dependency graphs and minimaps to simplify understanding.
Recent work by McNutt et al. mcnutt2023design explores how AI-powered code assistants can be optimized for computational notebook environments that have unique demands, which includes enhancing design features in support of iterative, exploratory coding. This points out the challenges in integrating AI-driven code suggestions into the typical nonlinear workflow inherent to notebooks and calls for meaningful customizability of design features.
Venkatesh et al. venkatesh2023enhancing introduces a tool-based approach (HeaderGen) to improve notebook comprehensibility through static analysis and automated header generation, our work focuses on understanding and predicting CU by analyzing code metrics. HeaderGen primarily enhances navigation and structural coherence in machine learning notebooks by classifying functions and adding markdown headers, which aids in comprehension.
The studies mentioned attempted to measure code quality using various features derived from Jupyter notebooks, but none of them measured the extent to which each feature can affect CU. Some of these studies only analyzed a small number of notebooks, which may not be representative of the broader population of Jupyter notebooks on GitHub. There is currently no widely accepted set of metrics for evaluating Jupyter notebook code quality. Different studies may use different metrics, making it difficult to compare results across studies. Our study is different and complements these studies by applying a machine learning model to determine if each metric is can affect CU. We also report which features have a significant impact on code comprehension and which features have an insignificant effect.
8.2 Measuring CU Based on Opinion
@@
Scalabrino et al. scalabrino2019automatically conducted an extensive evaluation of 121 existing, as well as new, code-related, documentation-related, and developer-related metrics. They attempted to correlate each metric with understandability and build models that combine metrics to assess understandability. To do this, they used 444 human evaluations from 63 developers. One of the criteria they used to measure CU is called Perceived Binary Understandability (PBU), which is based on developers’ opinions. This is a binary categorical variable that is true if a developer perceives that they understood a given code, and false otherwise. Finally, they obtained a bold negative result: none of the 121 experimented metrics is able to capture CU. As in the previous article, the methodology of our study is based on collecting opinions from a limited number of developers. The research scope of this article was limited to Java-based software.
Lavazza et al. lavazza2023empirical examined software understandability using metrics like Lines of Code, Cyclomatic Complexity, and Cognitive Complexity [^8] Cognitive Complexity measures how difficult code is to understand by evaluating its control flow and nesting. It penalizes deep nesting and complex logic while rewarding readability-friendly constructs, focusing on how humans perceive code. campbell2018cognitive . Their findings suggest that Cognitive Complexity does not significantly outperform traditional metrics in predicting CU, emphasizing the role of human factors alongside structural characteristics. They measured understandability through task completion times and applied machine learning to explore correlations between code measures and CU.
@@
The main challenge posed by the approaches presented in these studies is that they solely rely on user votes on software repositories. These metrics may be affected by factors beyond code comprehensibility, including ease of use, algorithm accuracy, popularity, and marketing considerations.
Despite the growing adoption of computational notebooks in data science and research, existing studies have primarily focused on notebook quality, reproducibility, and usability, without specifically addressing automated code understandability (CU) classification using notebook metrics. To the best of our knowledge, no prior work has systematically examined CU classification using a comprehensive set of 34 notebook metrics as classification features. Furthermore, no existing study has introduced an automated labeling approach like UOCU to assess CU. While some studies have explored aspects of notebook quality, these works have not proposed a structured methodology for CU classification, nor have they developed datasets and models tailored for this task. Our study fills this gap by presenting a novel approach that leverages notebook metrics for CU classification, providing a foundation for further exploration in this domain.
@@ 9.1 Internal Validity
It is a limitation of our approach that it only considers notebooks with more than 500 views for model fine-tuning. This threshold ensures a sufficient number of user comments for analysis but may introduce selection bias by excluding notebooks with fewer views. Consequently, our findings may not fully generalize to less-viewed notebooks, which could exhibit different understandability characteristics.
@@ 9.2 External Validity
While our study focused on the metrics and features that could be feasibly derived from the Kaggle platform, it provides a concrete example for future studies to build upon and further extend these criteria.
Another limitation of our study is that our classification model categorizes notebooks into only two groups: normal and good CU. While this simplifies the analysis and ensures robust classification, it does not account for notebooks with poor CU or provide a more fine-grained ranking. Future work could explore multi-class or regression-based approaches to provide a more nuanced assessment of CU.
@@ 9.3 Construct Validity
Construct validity refers to whether the theoretical constructs in the study are accurately measured easterbrook2008selecting . In our study, a potential threat to construct validity is the use of upvotes as a proxy for measuring CU. While upvotes have been used in previous studies as proxies for code quality and user satisfaction, they may not always reflect true understandability, as they can be influenced by factors like the notebook’s popularity or visibility.
To address this, we incorporated not only upvotes but also user comments and view counts into our analysis. By factoring in user comments, we ensured that feedback related to understandability was directly considered. Additionally, by filtering out notebooks with high view counts but relatively low upvotes, we reduced the risk of including notebooks that gained upvotes merely due to exposure rather than understandability. This multifaceted approach helps mitigate the limitations of upvotes as a sole measure of CU.
Furthermore, our large dataset and the use of additional baselines, such as markdown usage and code complexity, contribute to a more comprehensive and reliable assessment of CU.
@@ 9.4 Replicability
We have made the experimental setup, data, and source code fully accessible to ensure the replicability of our results. The complete source code for our experiments, is publicly available at https://github.com/ISE-Research/NotebookCU. This repository provides detailed instructions for setting up the environment and reproducing the outcomes presented in this paper, enabling fellow researchers to validate and extend our findings. Additionally, our code, dataset, and models are available at https://doi.org/10.5281/zenodo.14827719.
@@ 10 Conclusions and Future Work
COMPLIANCE WITH ETHICAL STANDARDS
Funding
This research received no external funding.
Ethical Approval
Not applicable. This study does not involve human participants or animal research requiring ethical approval.
Informed Consent
Not applicable. No human subjects were directly involved in experiments requiring informed consent.
Data Availability Statement
@@
Conflict of Interest
@@
Appendix A - Performance Tier Calculation
Appendix B - Annotation Guidelines for Classifying User Comments
Annotators should classify each comment according to its relevance to the understandability of the notebook. The comments fall into one of three categories:
  1. 1.
    Positive in Terms of Understandability A comment should be classified as positive if it expresses appreciation, praise, or specific feedback regarding the clarity, ease of understanding, or overall readability of the notebook. Positive comments often contain words like “clear”, “understandable”, “well-explained” or “easy to follow”.
    Examples:
    • “Great job explaining the steps; it made the analysis very easy to understand!”
    • “This notebook is incredibly clear and easy to follow. Thanks for sharing!”
  2. 2.
    Negative in Terms of Understandability A comment should be classified as negative if it highlights issues with the notebook’s clarity, difficulty in understanding, or areas where explanations are confusing or lacking. Negative comments may include phrases such as “hard to understand”, “confusing”, or “lacks explanation”.
    Examples:
    • “It’s difficult to follow the logic in this notebook.”
    • “The code is unclear, and there’s not enough explanation.”
  3. 3.
    No Relation to Understandability A comment should be classified as no relation to understandability if it does not directly address the notebook’s clarity or ease of understanding. This category includes comments that are unrelated to understandability, such as compliments on code efficiency, suggestions for additional features, general appreciation, or social interactions.
    Examples:
    • “Good job on the data preprocessing steps!”
    • “Thanks for sharing!”
    • “Could you add more visualizations?”
Appendix C - Distribution of Comments
@@
Table 12 contains comments from Kaggle users along with labels indicating the impact of these comments on Code Understandability):
  • Kaggle_Id: Unique identifier for a Kaggle comment written by the user.
  • Comment: The text of the comment provided by the user.
  • Positive CU: Indicates whether the comment positively contributes to the understanding of the code. A ’1’ means the comment is seen as helpful for improving code understandability; otherwise, it is left blank.
  • CU-unrelated: Indicates whether the comment is unrelated to the code understandability. A ’1’ means the comment does not pertain to the code understandability; otherwise, it is left blank.
  • Negative CU: Indicates whether the comment negatively impacts the understanding of the code. A ’1’ means the comment is viewed as detrimental to code understandability; otherwise, it is left blank.
Appendix D - Human Labeling Guideline for Notebook Classification
Purpose
The goal of this human labeling task is to categorize notebooks into two categories: “Normal Code Understandability” (NCU) and “Good Code Understandability” (GCU) (Our initial observation on user comments cause to remove any label related to “Weak Code Understandability”). This task is designed to assess the performance of the CU scores introduced in our study.
Definitions
  • Code Understandability (CU): Code Understandability, also referred to as program understandability or code comprehensibility, is the active process by which developers acquire knowledge about a software system through exploring software artifacts, reading source code, and consulting documentation [17].
  • Normal Code Understandability (NCU): Notebooks labeled as NCU are functional and meet basic standards of readability but lack advanced organizational features, detailed explanations, or may require additional effort to fully grasp the workflow. They may contain redundant code, inconsistent formatting, or limited documentation.
  • Good Code Understandability (GCU): Notebooks classified as GCU exhibit exceptional clarity, logical flow, thorough documentation, well-structured code cells, and effective use of visualizations. These notebooks require little effort for one to comprehend.
Characteristics to Consider
According to scientific sources, the normal conditions for understanding notebook code include the following characteristics:
  1. 1.
    Logical Structure and Flow: Ensures that the content is organized sequentially and coherently [9].
  2. 2.
    Descriptive Code and Documentation: Provides summaries to explain the code’s purpose, functionality, and insights [37].
  3. 3.
    Error Handling and Edge Cases: Handles errors and edge cases gracefully [52].
The assigned notebooks should evaluated independently by scoring them based on these characteristics. The scoring process is conducted using a numerical scale where a higher score reflects better understandability. The individual scores are then averaged, and a threshold-based approach is employed: If the average score exceeds half of the maximum possible score, the notebook is classified as GCU; otherwise, it is labeled as NCU. The final classification follows a binary scheme (GCU/NCU) to simplify interpretation and provide actionable insights.
For robustness, each notebook is reviewed by three independent annotators, and a consensus approach is followed to ensure reliability. A notebook is classified as GCU only if at least two out of three annotators agree on its classification. In cases of disagreement, additional review steps are taken to refine the assessment.
Examples
The following are examples of notebooks in each class:
Examples of NCU:
  • https://www.kaggle.com/code/scirpus/standard-gp/notebook
  • https://www.kaggle.com/code/alexandreh13/lung-cancer-simple-analysis-and-classification/notebook
  • https://www.kaggle.com/code/mayer79/m5-forecast-dept-by-dept-and-step-by-step/notebook
Examples of GCU:
  • https://www.kaggle.com/code/junkal/breast-cancer-prediction-using-machine-learning/notebook
  • https://www.kaggle.com/code/jneupane12/analysis-of-movielens-dataset-beginner-sanalysis/notebook
  • https://www.kaggle.com/code/sathishrao/iris-data-set/notebook
Instructions for Annotators
To complete the annotation task, please follow these steps:
  1. 1.
    Access the Google Form link provided to you.
  2. 2.
    Click on each notebook link within the form.
  3. 3.
    Carefully review each notebook, considering the definitions and characteristics mentioned above.
  4. 4.
    Score the notebook for each characteristic based on a numerical scale.
  5. 5.
    If the average score across all characteristics exceeds half of the maximum possible score, label the notebook as “GCU” (Good Code Understandability); otherwise, label it as “NCU” (Normal Code Understandability.
  6. 6.
    Assign a label to each notebook, as shown in Figure 5: either NCU or GCU as per criteria.
  7. 7.
    Submit your responses through the form upon completing the evaluation.
Figure 5: A sample for classify notebooks